The City of Decatur finds that salaries for residents working in the technology sector follow an approximately normal distribution with a mean of $42,800 and a standard deviation of $8,365. To the nearest hundredth, what Z-Score would be used to determine the percentage of residents working in the technology sector who earn more than $30,000?
Answer #1

Solution :

Given that,

mean = = 42800

standard deviation = =8365

P(x >30000 ) = 1 - P(x< 30000)

= 1 - P[(x -) / < (30000-42800) / 8365]

= 1 - P(z <-1.53 )

Using z table

= 1 - 0.0630

= 0.9370

=93.70%